# Jarník-Besicovitch type theorems for semisimple algebraic groups

### v1

@@ 4. A lower bound for the Hausdorff dimension of $S_{\gamma}^{c}$
where $\tilde{K}_{2}=B_{\tilde{\mathbf{H}}_{1}}(\theta_{2})\cdot K_{2}$ and $B_{\tilde{\mathbf{H}}_{1}}(\theta_{1},T)=a_{-T}\cdot B_{\tilde{\mathbf{H}}_{e}}(\theta_{1})\cdot a_{T}$. Here the constants $\theta_{1}$ and $\theta_{2}$ depend only on $K_{i}$’s, $\tilde{\mathbf{G}}$ and $\tilde{\Gamma}$.
Lemma 5.5.
We have
$\max\{\alpha(a_{1}):\alpha\in\Phi(\tilde{\mathbf{F}}_{e})\}>0.$
Consequently, using the same argument as in §4, we can conclude that
$\dim_{H}(S_{\tau}^{c}\cap\tilde{\mathbf{G}}(\mathbb{R}))\geq\dim\tilde{\mathbf{G}}-\sum_{\alpha\in\Phi(\tilde{\mathbf{F}}_{e})}\frac{\frac{\alpha(a_{1})}{\beta_{0}(a_{-1})-\tau}-\frac{\alpha(a_{1})}{\beta_{0}(a_{-1})}}{\frac{\tilde{\nu}_{0}(a_{1})}{\beta_{0}(a_{-1})-\tau}}$
where $\tilde{\nu}_{0}$ is the $\mathbb{Q}$-root in $R_{u}(\overline{\tilde{\mathbf{P}}}_{0})$ such that
@@
Note that $\mathbf{G}_{i}(\mathbb{R})$ $(s+1\leq i\leq k)$ all commute with $\{a_{t}\}_{t\in\mathbb{R}}$, and an element $g\in S_{\tau}^{c}$ if and only if $h\cdot g\in S_{\tau}^{c}$ for any $h\in\mathbf{G}_{i}(\mathbb{R})$ $(s+1\leq i\leq k)$. Therefore, we have
$\dim_{H}(S_{\tau}^{c})=$ $\dim_{H}(S_{\tau}^{c}\cap\tilde{\mathbf{G}}(\mathbb{R}))+\sum_{i=s+1}^{k}\dim\mathbf{G}_{i}(\mathbb{R})$
$\geq$ $\dim\mathbf{G}-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\frac{\frac{\alpha(a_{1})}{\beta_{0}(a_{-1})-\tau}-\frac{\alpha(a_{1})}{\beta_{0}(a_{-1})}}{\frac{\nu_{0}(a_{1})}{\beta_{0}(a_{-1})-\tau}}$
@@
$\dim_{H}(S_{\gamma}^{c}\cap U)\geq\dim R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\frac{\frac{\alpha(a_{1})}{\beta_{0}(a_{-1})-\tau}-\frac{\alpha(a_{1})}{\beta_{0}(a_{-1})}}{\frac{\nu_{0}(a_{1})}{\beta_{0}(a_{-1})-\tau}}.$
@@
Note that $\mathbf{P}_{0}\cdot R_{u}(\overline{\mathbf{P}}_{0})$ is a Zariski open dense subset in $\mathbf{G}$, and if $x\in S_{\gamma}^{c}\cap U$, then for any $g\in\mathbf{P}_{0}(\mathbb{R})$, $g\cdot x$ is also an element in $S_{\gamma}^{c}$. This implies that for any open bounded subset $\tilde{U}\subset\mathbf{G}(\mathbb{R})$
$\dim_{H}(S_{\gamma}^{c}\cap\tilde{U})\geq\dim\mathbf{G}-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\frac{\frac{\alpha(a_{1})}{\beta_{0}(a_{-1})-\tau}-\frac{\alpha(a_{1})}{\beta_{0}(a_{-1})}}{\frac{\nu_{0}(a_{1})}{\beta_{0}(a_{-1})-\tau}}.$
6. An upper bound for the Hausdorff dimension of $S_{\gamma}^{c}$
@@
In this section, we prove Theorem 1.4. We will compute an upper bound of $\dim_{H}S_{\gamma}^{c}$ by constructing open covers of the subset $S_{\gamma}^{c}$. Recall that
$\mathcal{C}_{\bar{I}}:=\bigcup_{i\in\bar{I}}w_{i}\mathbf{F}_{w_{i}}(\mathbb{R}).$
@@
for some $g\in\mathbf{G}(\mathbb{R})$ and $r>0$. Then there exist $\tilde{g}\in\mathcal{C}_{\bar{I}}$ and a discrete primitive subgroup
@@
for some Siegel set $S_{\eta,\Omega}=K\cdot T_{\eta}\cdot\Omega$ and some finite subset $\mathcal{K}\subset\mathbf{G}(\mathbb{Q})$ where
@@
for some $\tilde{k}$ in a fixed compact subset in $\mathbf{G}(\mathbb{R})$.
@@
$\|\rho(a\cdot x)\cdot y\|\sim\|\rho(\tilde{k}\cdot a\cdot x)\cdot y\|\leq r.$
@@
according to the decomposition of the complex vector space $V=\oplus_{\beta}V_{\beta}$ in $\rho$ into weight spaces $V_{\beta}$ relative to $\mathbf{T}$, where $y_{\beta}\in V_{\beta}$. Let $\tilde{\beta}$ be a weight among the weights $\beta$’s in the summation such that $y_{\tilde{\beta}}\neq 0$. Then by the structure of the representation $\rho$, we compute that
@@
Note that $y\in\mathbb{Z}^{d}$ and $x\in\mathcal{K}$, and there is a positive lower bound for $\|y_{\tilde{\beta}}\|$ depending only on $\rho$ and $\mathcal{K}$. On the other hand, there exists a unique discrete primitive subgroup $\tilde{\Lambda}\subset\mathbb{Z}^{d}$ such that $\rho(x)\cdot\tilde{\Lambda}$ is a discrete and Zariski-dense subgroup in $V_{\beta_{0}}$, and for any $w\in\tilde{\Lambda}\setminus\{0\}$ we have
(2) $\|\rho(\tilde{k}\cdot a\cdot x)\cdot w\|\sim\|\rho(a)\cdot\rho(x)\cdot w\|=e^{\beta_{0}(a)}\|\rho(x)\cdot w\|.$
Comparing equations (1) and (2) and using the fact that $x\in\mathcal{K}\subset\mathbf{G}(\mathbb{Q})$, $a\in T_{\eta}$ and the relation between $\tilde{\beta}$ and $\beta_{0}$ according to the structure of the representation $\rho$, one can deduce that
$e^{\beta_{0}(a)}\lesssim_{\eta}e^{\tilde{\beta}(a)}\lesssim r$
and for any $w\in\tilde{\Lambda}\setminus\{0\}$
@@
Let $\Lambda_{g}=\rho(g)(\rho(\gamma^{-1})\tilde{\Lambda})$. Then we have
@@
By the Bruhat decomposition, we know that
@@
$\tilde{k}^{-1}\in\bigcup_{i\in\bar{I}}\mathbf{P}_{\beta_{0}}(\mathbb{R})\cdot w_{i}\cdot\mathbf{F}_{w_{i}}(\mathbb{R}).$
By the fact that $\mathbf{P}_{\beta_{0}}$ stabilizes $V_{\beta_{0}}$, we conclude that there exists $\tilde{g}\in\mathcal{C}_{\bar{I}}$ such that
$\tilde{k}^{-1}\in\mathbf{P}_{\beta_{0}}(\mathbb{R})\cdot\tilde{g}\textup{ and }\rho(\tilde{g})\cdot\Lambda_{g}\subset V_{\beta_{0}}.$
This completes the proof of the proposition. ∎
@@
Let $\epsilon_{0}>0$ which we will determine later. Let $g$ be an element in $S_{\tau}^{c}\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$. Then by definition, there exists a sequence $\{t_{k}\}\subset\mathbb{N}$ such that
$\delta(\rho(a_{t_{k}}g)\cdot\mathbb{Z}^{d})\leq\epsilon_{0}\cdot e^{-\tau t_{k}}.$
@@
By Proposition 6.1, for each $k\in\mathbb{N}$, there exist $\tilde{g}_{k}\in\mathcal{C}_{\bar{I}}$ and a discrete subgroup $\Lambda_{k}\subset\rho(a_{t_{k}}g)\cdot\mathbb{Z}^{d}$ such that
@@
$\operatorname{Vol}(\Lambda_{k})\lesssim(\epsilon_{0}e^{-\tau t_{k}})^{\dim V_{\beta_{0}}}\textup{ and }\rho(\tilde{g}_{k})\cdot\Lambda_{k}\textup{ is Zariski dense in }V_{\beta_{0}}.$
@@
Let $\tilde{g}_{k}=w_{i,k}\cdot f_{i,k}\in\mathcal{C}_{\bar{I}}$ where $w_{i,k}\in\{w_{i}\}_{i\in\bar{I}}$ and $f_{i,k}\in\mathbf{F}_{w_{i,k}}(\mathbb{R})$. Then it implies that
$\tilde{q}_{k}:=w_{i,k}\cdot f_{i,k}\cdot a_{t_{k}}\cdot g=w_{i,k}\cdot a_{t_{k}}\cdot(a_{-t_{k}}f_{i,k}a_{t_{k}})\cdot g$
$q_{k}:=(w_{i,k}a_{-t_{k}}w_{i,k}^{-1})\cdot\tilde{q}_{k}=w_{i,k}(a_{-t_{k}}f_{i,k}a_{t_{k}})g$
are rational elements in $\mathbf{G}(\mathbb{R})$. In the $\dim V_{\beta_{0}}$-exterior product space of $V$, we can compute that
$d(q_{k})=e^{\beta_{0}(w_{i,k}a_{-t_{k}}w_{i,k}^{-1})\cdot\dim V_{\beta_{0}}}\cdot d(\tilde{q}_{k})$
$d(\tilde{q}_{k})\cdot e_{V_{\beta_{0}}}=\rho_{\beta_{0}}(\tilde{g}_{k})\cdot\left(\bigwedge^{\dim V_{\beta_{0}}}\Lambda_{k}\right),\quad\rho_{\beta_{0}}(\tilde{g}_{k}^{-1})\cdot e_{V_{\beta_{0}}}=\left(\bigwedge^{\dim V_{\beta_{0}}}\Lambda_{k}\right)/d(\tilde{q}_{k})$
$\|\rho_{\beta_{0}}(\tilde{g}_{k}^{-1})\cdot e_{V_{\beta_{0}}}\|=$ $\operatorname{Vol}(\Lambda_{k})/d(\tilde{q}_{k})$
$\lesssim$ $(\epsilon_{0}\cdot e^{-\tau t_{k}})^{\dim V_{\beta_{0}}}\cdot e^{\beta_{0}(w_{i,k}a_{-t_{k}}w_{i,k}^{-1})\cdot\dim V_{\beta_{0}}}/d(q_{k})$
@@
where $e_{V_{\beta_{0}}}=\bigwedge^{\dim V_{\beta_{0}}}V_{\beta_{0}}$ is the unit vector which represents the vector space $V_{\beta_{0}}$ in the $\dim V_{\beta_{0}}$-exterior product space of $V$, and $\bigwedge^{\dim V_{\beta_{0}}}\Lambda_{k}$ is the vector constructed from a $\mathbb{Z}$-basis of $\Lambda_{k}$ and represents the lattice $\Lambda_{k}$. The implicit constants here depend only on $\mathbf{G}$, $\Gamma$ and $\rho$.
@@
Fix an arbitrary $i\in\bar{I}$. For any $R>0$, define
$E_{w_{i}}(R):=\{f\in\mathbf{F}_{w_{i}}(\mathbb{R}):\|\rho_{\beta_{0}}(w_{i}fw_{i}^{-1})\cdot e_{V_{\beta_{0}}}\|\leq R\}.$
@@
$\Psi_{w_{i}}:\mathbf{F}_{w_{i}}(\mathbb{R})\to\bigwedge_{j=1}^{\dim V_{\beta_{0}}}V,\quad\Psi_{w_{i}}(x)=\rho_{\beta_{0}}(w_{i}xw_{i}^{-1})\cdot e_{V_{\beta_{0}}}.$
Note that $w_{i}\mathbf{F}_{w_{i}}w_{i}^{-1}\subset R_{u}(\overline{\mathbf{P}}_{\beta_{0}})$. Hence $\Psi_{w_{i}}$ is an isomorphism onto its image.

### v2

@@ 3. Counting rational elements
where $\tilde{K}_{2}=B_{\tilde{\mathbf{H}}_{e}}(\theta_{2})\cdot K_{2}$ and $B_{\tilde{\mathbf{H}}_{e}}(\theta_{1},T)=a_{-T}\cdot B_{\tilde{\mathbf{H}}_{e}}(\theta_{1})\cdot a_{T}$. Here the constants $\theta_{1}$ and $\theta_{2}$ depend only on $K_{i}$’s, $\tilde{\mathbf{G}}$ and $\tilde{\Gamma}$.
Consequently, using the same argument as in §4, we can conclude that for any $0\leq\tau(\psi)<\beta_{0}(a_{-1})$
$\dim_{H}(S_{\rho}(\psi)^{c}\cap\tilde{\mathbf{G}}(\mathbb{R}))\geq\dim\tilde{\mathbf{G}}-\sum_{\alpha\in\Phi(\tilde{\mathbf{F}}_{e})}\frac{\alpha(a_{1})\tau(\psi)}{\beta_{0}(a_{-1})\tilde{\nu}_{0}(a_{1})}$
where $\tilde{\nu}_{0}$ is any $\mathbb{Q}$-root in $R_{u}(\overline{\tilde{\mathbf{P}}}_{0})$ such that
@@
Note that $\mathbf{G}_{i}(\mathbb{R})$ $(s+1\leq i\leq k)$ commutes with $\{a_{t}\}_{t\in\mathbb{R}}$, and an element $g\in S_{\rho}(\psi)^{c}$ if and only if $h\cdot g\in S_{\rho}(\psi)^{c}$ for any $h\in\mathbf{G}_{i}(\mathbb{R})$ $(s+1\leq i\leq k)$. Therefore, we have
$\dim_{H}(S_{\rho}(\psi)^{c})=$ $\dim_{H}(S_{\rho}(\psi)^{c}\cap\tilde{\mathbf{G}}(\mathbb{R}))+\sum_{i=s+1}^{k}\dim\mathbf{G}_{i}(\mathbb{R})$
$\geq$ $\dim\mathbf{G}-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\frac{\alpha(a_{1})\tau(\psi)}{\nu_{0}(a_{1})\beta_{0}(a_{-1})}$
@@
$\dim_{H}(S_{\rho}(\psi)^{c}\cap U)\geq\dim R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\frac{\alpha(a_{1})\tau(\psi)}{\nu_{0}(a_{1})\beta_{0}(a_{-1})}.$
@@
Note that $\mathbf{P}_{0}\cdot R_{u}(\overline{\mathbf{P}}_{0})$ is a Zariski open dense subset in $\mathbf{G}$, and for any open bounded subset $\tilde{U}\subset\mathbf{G}(\mathbb{R})$, one can find open subsets $U_{1}$ and $U_{2}$ in $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$ and $\mathbf{P}_{0}(\mathbb{R})$ respectively such that
$U_{2}\cdot U_{1}\subset\tilde{U}$.
Moreover, by definition, if $x\in S_{\rho}(\psi)^{c}$, then for any $g\in\mathbf{P}_{0}(\mathbb{R})$, $g\cdot x$ is also an element in $S_{\rho}(\psi)^{c}$. This implies that
$\dim_{H}(S_{\rho}(\psi)^{c}\cap\tilde{U})\geq\dim U_{2}+\dim_{H}(S_{\rho}(\psi)^{c}\cap U_{1})\geq\dim\mathbf{G}-\sum_{\alpha\in\Phi(\mathbf{F}_{e})}\frac{\alpha(a_{1})\tau(\psi)}{\beta_{0}(a_{-1})\nu_{0}(a_{1})}.$
Finally, we study the case $\tau(\psi)=\beta_{0}(a_{-1})$ and conclude the proof of Theorem 1.4.
Proposition 5.6.
Let $\mathbf{G}$ be a connected semisimple algebraic group, $\rho$ a $\mathbb{Q}$-rational irreducible representation of $\mathbf{G}$, $\{a_{t}\}_{t\in\mathbb{R}}$ a one-parameter subgroup in $\mathbf{T}(\mathbb{R})$ and $\beta_{0}$ the highest weight defined as in §2. Let $\psi:\mathbb{R}_{+}\to\mathbb{R}_{+}$ with $\tau(\psi)=\beta_{0}(a_{-1})$. Then
  1. (1)
    If $\psi(t)\cdot e^{\beta_{0}(a_{-1})t}$ is bounded, then $S_{\rho}(\psi)^{c}=\emptyset$.
  2. (2)
    If $\psi(t)\cdot e^{\beta_{0}(a_{-1})t}$ is unbounded, then $S_{\rho}(\psi)^{c}\neq\emptyset$ and
    $\dim_{H}S_{\rho}(\psi)^{c}\geq\dim\mathbf{G}-\frac{\tau(\psi)}{\beta_{0}(a_{-1})\nu_{0}(a_{1})}\cdot\sum_{\alpha\in\Phi(R_{u}(\overline{\mathbf{P}}_{\beta_{0}}))}\alpha(a_{1}).$
Proof.
Suppose that $\psi(t)\cdot e^{\beta_{0}(a_{-1})t}$ is bounded. Then there exists $C>0$ such that
$\psi(t)\leq C\cdot e^{-\beta_{0}(a_{-1})t}\quad(t>0).$
Let $g\in\mathbf{G}(\mathbb{R})$. Since $\beta_{0}(a_{-1})>0$ is the fastest contracting rate of the $\{a_{t}\}_{t\in\mathbb{R}}$-action on $V$, we have
$\|\rho(a_{t})v\|\geq e^{-\beta_{0}(a_{-1})t}\|v\|\quad(\forall v\in\rho(g)\cdot\mathbb{Z}^{d}\setminus\{0\})$
and
$\delta(\rho(a_{t}\cdot g)\cdot\mathbb{Z}^{d})\geq e^{-\beta_{0}(a_{-1})t}\cdot\delta(\rho(g)\cdot\mathbb{Z}^{d}).$
One can find a constant $C_{g}>0$ depending on $g$ such that
$\delta(\rho(a_{t}\cdot g)\cdot\mathbb{Z}^{d})\geq C_{g}\cdot\psi(t)\quad(t>0)$
which shows that $g\in S_{\rho}(\psi)$ and $S_{\rho}(\psi)=\mathbf{G}(\mathbb{R})$. This proves the first claim.
Now suppose that $\psi(t)\cdot e^{\beta_{0}(a_{-1})t}$ is unbounded. We first consider the case where $\{a_{t}\}_{t\in\mathbb{R}}$ on $\mathbf{G}(\mathbb{R})^{0}/\Gamma$ is mixing. We may write
$\psi(t)\cdot e^{\beta_{0}(a_{-1})t}=e^{h(t)}$
where $h(t)$ is not bounded from above. Since $\tau(\psi)=\beta_{0}(a_{-1})$, we have
$\limsup_{t\to\infty}\frac{h(t)}{t}=0.$
Let $\{s_{k}\}_{k\in\mathbb{N}}$ be a divergent sequence in $\mathbb{R}_{+}$ such that $\lim_{k\to\infty}h(s_{k})=\infty.$ Then
$\lim_{k\to\infty}\frac{h(s_{k})}{s_{k}}=0.$
In §4, we construct a Cantor-type subset $\mathbf{A}_{\infty}$ in $R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$ with parameters $l_{k}$ and $t_{k}$ $(k\in\mathbb{N})$. Here to prove the proposition with $\tau(\psi)=\beta_{0}(a_{-1})$, we may take $l_{k}=e^{h(s_{k})/2}$, $t_{k}=s_{k}$ and $\tau_{k}=\beta_{0}(a_{-1})-h(s_{k})/(2s_{k})$ $(k\in\mathbb{N})$. Then
$t_{k}=\frac{\ln l_{k}}{\beta_{0}(a_{-1})-\tau_{k}}.$
Using the same argument in Lemma 4.3, one can prove that for any $p\in\mathbf{A}_{\infty}$
$\delta(\rho(a_{t_{k}}\cdot p)\mathbb{Z}^{d})$ $\lesssim e^{-\tau_{k}\cdot t_{k}}=e^{-\beta_{0}(a_{-1})t_{k}+h(t_{k})/2}=\psi(t_{k})\cdot e^{-h(t_{k})/2}\quad(k\in\mathbb{N})$
and $\mathbf{A}_{\infty}\subset S_{\rho}(\psi)^{c}\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$. By Theorem 4.2, we may estimate the Hausdorff dimension of $\mathbf{A}_{\infty}$ and obtain
$\dim_{H}S_{\rho}(\psi)^{c}\geq\dim\mathbf{G}-\frac{\tau(\psi)}{\beta_{0}(a_{-1})\nu_{0}(a_{1})}\cdot\sum_{\alpha\in\Phi(R_{u}(\overline{\mathbf{P}}_{\beta_{0}}))}\alpha(a_{1}).$
If the action of $\{a_{t}\}_{t\in\mathbb{R}}$ on $\mathbf{G}(\mathbb{R})^{0}/\Gamma$ is not mixing, then we may repeat the arguments in this section and reduce the problem in the homogeneous space $\tilde{\mathbf{G}}(\mathbb{R})^{0}/\tilde{\Gamma}$. This completes the proof of the proposition. ∎
6. An upper bound for the Hausdorff dimension of $S_{\rho}(\psi)^{c}$
@@
In this section, we prove Theorem 1.5. We will first compute an upper bound for the Hausdorff dimension of $S_{\rho}(\psi)^{c}\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$ by constructing open covers of the subset $S_{\rho}(\psi)^{c}\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$, and then derive Theorem 1.5 from the upper bound of $\dim_{H}S_{\rho}(\psi)^{c}\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$. We denote by
$\mathcal{C}:=\bigcup_{w\in{{}_{\mathbb{Q}}\overline{\mathcal{W}}}}w\mathbf{F}_{w}(\mathbb{R}).$
@@
for some $g\in\mathbf{G}(\mathbb{R})$ and $r>0$. Then there exist $\tilde{g}\in\mathcal{C}$ and a discrete primitive subgroup of rank $\dim V_{\beta_{0}}$
@@
for some Siegel set $S_{\eta,\Omega}=K\cdot T_{\eta}\cdot\Omega$ and some finite subset $\mathcal{K}\subset\mathbf{G}(\mathbb{Q})$, where
@@
for $\tilde{k}=k\cdot(a\cdot u\cdot a^{-1})$ in a fixed compact subset in $\mathbf{G}(\mathbb{R})$.
@@
$\|\rho(a\cdot x)\cdot y\|\asymp\|\rho(\tilde{k}\cdot a\cdot x)\cdot y\|\leq r.$
@@
according to the decomposition of $V=\oplus_{\beta}V_{\beta}$ in $\rho$ into the weight spaces $V_{\beta}$ relative to $\mathbf{T}$, where $y_{\beta}\in V_{\beta}$. Let $\tilde{\beta}$ be a weight among the weights $\beta$’s in the summation above such that $y_{\tilde{\beta}}\neq 0$. Then
@@
Note that $y\in\mathbb{Z}^{d}$ and $x\in\mathcal{K}\subset\mathbf{G}(\mathbb{Q})$, and there is a positive lower bound for $\|y_{\tilde{\beta}}\|$ depending only on $\rho$ and $\mathcal{K}$. So by equation (2), we have
$e^{\tilde{\beta}(a)}\lesssim r.$
On the other hand, there exists a unique discrete primitive subgroup $\tilde{\Lambda}\subset\mathbb{Z}^{d}$ such that $\rho(x)\cdot\tilde{\Lambda}$ is a discrete and Zariski-dense subgroup in $V_{\beta_{0}}$, and for any $w\in\tilde{\Lambda}\setminus\{0\}$ we have
(3) $\|\rho(\tilde{k}\cdot a\cdot x)\cdot w\|\asymp\|\rho(a)\cdot\rho(x)\cdot w\|=e^{\beta_{0}(a)}\|\rho(x)\cdot w\|.$
By the condition $a\in T_{\eta}$ and the relation between $\tilde{\beta}$ and $\beta_{0}$ according to the structure of the representation $\rho$, one can compute that
$e^{\beta_{0}(a)}\lesssim_{\eta}e^{\tilde{\beta}(a)}\lesssim r.$
Consequently, from equation (3) one can deduce that for any $w\in\tilde{\Lambda}\setminus\{0\}$
@@
[…]
@@
We know that
@@
$\tilde{k}^{-1}\in\mathbf{G}(\mathbb{R})=\bigcup_{w\in{{}_{\mathbb{Q}}\overline{\mathcal{W}}}}\mathbf{P}_{\beta_{0}}(\mathbb{R})\cdot w\cdot\mathbf{F}_{w}(\mathbb{R}).$
So there exists $\tilde{g}\in\mathcal{C}$ such that
$\tilde{k}^{-1}\in\mathbf{P}_{\beta_{0}}(\mathbb{R})\cdot\tilde{g}$.
By the fact that $\mathbf{P}_{\beta_{0}}$ stabilizes $V_{\beta_{0}}$, we conclude that
$\rho(\tilde{g})\cdot\Lambda_{g}\subset V_{\beta_{0}}.$
Note that all the implicit constants above depend only on $\mathbf{G}$, $\Gamma$ and $\rho$. This completes the proof of the proposition. ∎
@@
Let $g$ be an element in $S_{\rho}(\psi)^{c}\cap R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$. Then from Definition 1.1, one can deduce that there exists a divergent sequence $\{t_{k}\}\subset\mathbb{R}_{+}$ such that
$\delta(\rho(a_{t_{k}}g)\cdot\mathbb{Z}^{d})\leq\psi(t_{k}).$
Let $\epsilon>0$ be a sufficiently small number. By the definition of $\tau=\tau(\psi)$, we have
$\psi(t)\leq C_{\epsilon}\cdot e^{-(\tau-\epsilon)t}\quad(\forall t>0)$
for some $C_{\epsilon}>0$ depending only on $\epsilon$. Hence
$\delta(\rho(a_{t_{k}}g)\cdot\mathbb{Z}^{d})\leq C_{\epsilon}\cdot e^{-(\tau-\epsilon)t_{k}}.$
@@
Here by adding a multiplicative constant in the inequality above, we may assume that $t_{k}\in\mathbb{N}$. By Proposition 6.1, for each $k\in\mathbb{N}$, there exist $\tilde{g}_{k}\in\mathcal{C}$ and a discrete subgroup $\Lambda_{k}\subset\rho(a_{t_{k}}g)\cdot\mathbb{Z}^{d}$ of rank $\dim V_{\beta_{0}}$ such that
@@
$\operatorname{Vol}(\Lambda_{k})\lesssim(C_{\epsilon}\cdot e^{-(\tau-\epsilon)t_{k}})^{\dim V_{\beta_{0}}}\textup{ and }\rho(\tilde{g}_{k})\cdot\Lambda_{k}\textup{ is Zariski dense in }V_{\beta_{0}}.$
@@
Since $_{\mathbb{Q}}\overline{\mathcal{W}}$} is finite, by passing to a subsequence, we may write $\tilde{g}_{k}=w\cdot f_{k}\in\mathcal{C}$ where $w\in{{}_{\mathbb{Q}}\overline{\mathcal{W}}}$ is fixed for all $k\in\mathbb{N}$ and $f_{k}\in\mathbf{F}_{w}(\mathbb{R})$. Then by definition
$\tilde{q}_{k}:=\tilde{g}_{k}\cdot a_{t_{k}}\cdot g=w\cdot f_{k}\cdot a_{t_{k}}\cdot g=w\cdot a_{t_{k}}\cdot(a_{-t_{k}}f_{k}a_{t_{k}})\cdot g$
$q_{k}:=(wa_{-t_{k}}w^{-1})\cdot\tilde{q}_{k}=w\cdot(a_{-t_{k}}f_{k}a_{t_{k}})\cdot g$
are rational elements in $\mathbf{G}(\mathbb{R})$.
@@
In the $\dim V_{\beta_{0}}$-exterior product space of $V$, let $e_{V_{\beta_{0}}}=e_{1}\wedge e_{2}\wedge\cdots\wedge e_{\dim V_{\beta_{0}}}$ be the unit vector which represents the vector space $V_{\beta_{0}}$ (see §1), and $\wedge^{\dim V_{\beta_{0}}}\Lambda_{k}$ the vector in $\bigwedge^{\dim V_{\beta_{0}}}V$ which is constructed from a $\mathbb{Z}$-basis of $\Lambda_{k}$ and represents the lattice $\Lambda_{k}$. Then we can compute that
$0pt(q_{k})=e^{\beta_{0}(wa_{-t_{k}}w^{-1})\cdot\dim V_{\beta_{0}}}\cdot 0pt(\tilde{q}_{k})$
$0pt(\tilde{q}_{k})\cdot e_{V_{\beta_{0}}}=\rho_{\beta_{0}}(\tilde{g}_{k})\cdot\left(\wedge^{\dim V_{\beta_{0}}}\Lambda_{k}\right),\quad\rho_{\beta_{0}}(\tilde{g}_{k}^{-1})\cdot e_{V_{\beta_{0}}}=\left(\wedge^{\dim V_{\beta_{0}}}\Lambda_{k}\right)/0pt(\tilde{q}_{k})$
$\|\rho_{\beta_{0}}(\tilde{g}_{k}^{-1})\cdot e_{V_{\beta_{0}}}\|=$ $\operatorname{Vol}(\Lambda_{k})/0pt(\tilde{q}_{k})$
$\lesssim$ $(C_{\epsilon}\cdot e^{-(\tau-\epsilon)t_{k}})^{\dim V_{\beta_{0}}}\cdot e^{\beta_{0}(wa_{-t_{k}}w^{-1})\cdot\dim V_{\beta_{0}}}/0pt(q_{k}).$
The implicit constants here depend only on $\mathbf{G}$, $\Gamma$ and $\rho$. For convenience, we denote by
$\lambda_{w}(\epsilon):=(\beta_{0}(wa_{-1}w^{-1})-(\tau-\epsilon))\cdot\dim V_{\beta_{0}}.$
@@
Let $w\in{{}_{\mathbb{Q}}\overline{\mathcal{W}}}$. For any $R>0$, define
$E_{w}(R):=\{f\in\mathbf{F}_{w}(\mathbb{R}):\|\rho_{\beta_{0}}(wfw^{-1})\cdot e_{V_{\beta_{0}}}\|\leq R\}.$
@@
$\Psi_{w}:\mathbf{F}_{w}(\mathbb{R})\to\bigwedge^{\dim V_{\beta_{0}}}V,\quad\Psi_{w}(x)=\rho_{\beta_{0}}(wxw^{-1})\cdot e_{V_{\beta_{0}}}.$
Note that $w\mathbf{F}_{w}w^{-1}\subset R_{u}(\overline{\mathbf{P}}_{\beta_{0}})$ and the stabilizer of $\mathbb{C}\cdot e_{V_{\beta_{0}}}$ is equal to $\mathbf{P}_{\beta_{0}}$. Hence $\Psi_{w}$ is an isomorphism onto its image.